I roll 5 dice. What is the probability of having a sum of 26 and that one of the dice is a 1?
Answer #1

The outcome of each of dice can be {1,2,3,4,5,6}
If Xi be the outcome of ith dice, X1 + X2 + X3 + X4 + X5 = 26 but it is also given that one if dice is a 1.

which means, Sum of other four dice = 26-1 = 25. But maximum sum of four dice can be 24.

Hence required probability = 0